Find the real zeros of the function P(x)=3(x+8)^2-(x+8)+4
Answer by lwsshak3(11628)

Find the real zeros of the function
P(x)=3(x+8)^2-(x+8)+4
**
P(x)=3(x+8)^2-(x+8)+4=0
let u=(x+8)
3u^2-u+4=0
b^2-4ac=1-4*3*4<0
no real solutions